OpenStudy (anonymous):

The circumference of a standard baseball is about 9in. About how many square in. of horsehide are required to cover 100 balls, to the nearest whole number?

OpenStudy (noseboy908):

Firstly, you need to use the circumference formula to determine the radius. So, can you please state that?

OpenStudy (anonymous):

A=\[\pi r^2 \]

OpenStudy (noseboy908):

That's correct, but what is the formula for circumference?

OpenStudy (anonymous):

C=\[\pi d\]

OpenStudy (noseboy908):

Alright, so fill in the one thing you know in that formula: the total circumference. Then, restate the formula for me.

OpenStudy (anonymous):

9=\[\pi *100\]

OpenStudy (noseboy908):

Where did you get 100 for the diameter? Is that given to you?

OpenStudy (anonymous):

It said that how many square in. of horsehide are required to cover 100 balls.

OpenStudy (noseboy908):

You're right about that, but you put it in the place of the diameter. The number of total balls is used, but not for a few more steps. So, restate the formula as you just did, but replace the 100 you put in with "d" again.

OpenStudy (anonymous):

9=\[\pi *d\]

OpenStudy (noseboy908):

Correct, now, given that your answer will be a decimal in the end anyway, solve for d. Be as precise as possible with the pie.

OpenStudy (noseboy908):

(I say that because sometimes they want it in terms of pie.)

OpenStudy (anonymous):

28.27 or if I round it up it would be 28

OpenStudy (noseboy908):

Show me how you got that. Remember that in order to remove something being multiplied from one side, you must divide it from both sides.

OpenStudy (anonymous):

 I was confused at the end